OpenStudy (anonymous):

How many solutions does the equation have? 4(2x +3) = 2(3x - 4) A. One solution B. Infinite solutions C. No solution

OpenStudy (anonymous):

it is no solution

OpenStudy (anonymous):

Thank you so much.

OpenStudy (anonymous):

hey when ever u need help just tag me ok

OpenStudy (anonymous):

Can I give you a couple more?

OpenStudy (anonymous):

yea but i got to switch classes but i will be on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

a+5=1/5(15+5a)

OpenStudy (anonymous):

4x + 2(x-3) =8x + 12

OpenStudy (anonymous):

um is there answers '

OpenStudy (anonymous):

They're the same as the first one. They're asking about the solutions.

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

And thanks again. I have to get a whole bunch of work done before Easter or I'll fail my grade and I can't get my lip pierced unless I pass lol. So thanks a ton.

OpenStudy (anonymous):

don't worry about it that im tryna do get my ears pirced

OpenStudy (anonymous):

but I need to pass

OpenStudy (anonymous):

I got one ear pierced when I was like 10 but my school made me take it out. I'm probably going to start going to public school next year so they'll probably try and make me take my snake-bites out but I'm just gonna be rebel and.....rebel? xD

OpenStudy (anonymous):

lol alright for the first one u are gonna simplify

OpenStudy (anonymous):

then multiply the rest

OpenStudy (anonymous):

for the second one

OpenStudy (anonymous):

then for the 1st one u do the same

OpenStudy (anonymous):

What all do I simplify? Omg sorry I'm so dumb when it comes to math.

OpenStudy (anonymous):

hey me to but I get through it

OpenStudy (anonymous):

the top numbers

OpenStudy (anonymous):

In the fractions? Because there's only one fraction and its top number is 1 and I can't simplify 1.

OpenStudy (anonymous):

@ortiz09 ? I'm hella confused right now.

OpenStudy (anonymous):

You there? @ortiz09

OpenStudy (tkhunny):

For the very fist question in this thread that is too long, try x = -10. Let's work on actually helping, rather than either 1) Giving out answers or 2) Giving out incorrect answers.